A ladder is leaning against the side of a house in order to reach a window that is 25 feet off the ground. The angle formed between the ladder and the ground is 52 degrees. How far from the base of the house is the base of the ladder? Be sure to show all of your work. Round your answer to the nearest tenth of a foot.
Answer by TimothyLamb(4379)

---
tan(A) = opp/adj
tan(52) = h/b
tan(52) = 25/b
b = 25/tan(52)
b = 19.5 ft
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php